Hi, Please help me with this Algebra problem. Thank you in advance.
Solve using the multiplication principle. Also don't forget to check.
a/4=13

Answer by stanbon(57377)

a/4=13
---
Multiply both sides by 4 to get:
a = 4*13
a = 52
----
Check:
a/4 = 13
52/4 = 13
13 = 13
============
Cheers,
Stan H.
============